A rental agent claims that the mean monthly rent, , for apartments on the east side of town is less than $750 . A random sample of 16 monthly rents for apartments on the east side has a mean of $748 , with a standard deviation of $22. If we assume that the monthly rents for apartments on the east side are normally distributed, is there enough evidence to conclude, at the 0.05 level of significance, that is less than $750 ?
Answer by stanbon(75887)

A rental agent claims that the mean monthly rent, , for apartments on the east side of town is less than $750.
-------
A random sample of 16 monthly rents for apartments on the east side has a mean of $748 , with a standard deviation of $22.
-------
If we assume that the monthly rents for apartments on the east side are normally distributed, is there enough evidence to conclude, at the 0.05 level of significance, that is less than $750?
----------------------------
Ho: u >= 750
Ha: u < 750 (claim)
-------------
x-bar = 748
Test statistic: t(748) = (748-750)/[22/sqrt(16)] = -0.3636
----
p-value = P(t<-0.3535 when df = 15) = tcdf(-100,-0.3636,15) = 0.3606
----
Since the p-value is greater than 5%, fail to reject Ho at
the 5% level of significance.
==================================
The test results do not support the claim)
==================================
Cheers,
Stan H.
===========